OpenStudy (ucsdsongbird):

use Lagrange Multipliers to find max & min of f(x,y)=5xy constrained by 9x+8y=180 I already found the critical point (10, 45/4). f(10,45/4)=1125/2, this is the max value. the answer book says there is no minimum value, but how do you know? for example f(20,0)=0, why isn't this the min value?

OpenStudy (zarkon):

a simple way to see why it has no min is to look at the following \[9x+8y=180\Rightarrow8y=180-9x\] then \[f(x,y)=5xy=\frac{5}{8}x8y=\frac{5}{8}x(180-9x)\] which is a parabola which opens down. therefore it has no min.

OpenStudy (irishboy123):

the multiplier simplifies the algebra but in my experience is pretty bad at ID'ing the nature of the extrema turning it here into a single var problem makes that poss. a lot of the time, it turns on the physical nature of the function